Grade 12th passTrigonometry

In the properties of triangle if the triangle ABC has a=2√3 b= 4 then A=60 then find perimeter of the triangle

Anusha
7 Years agoGrade 12th pass

1 Answer

Arun
ApprovedApproved Tutor Answer7 Years ago
Dear student
 
by sin rule
 
a/ sinA = b/ sinB
 
sin B = b sin A/ a = 1
 
hence B = 90 degrees
 
hence it is a right angled triangle
 
c^2 + a^2 = b^2
 
c^2 + 12 = 16
c^2 = 4
c = 2
 
hence perimeter = a + b+ c = 6 + 2 sqrt (3)